# Hilbert-Kunz multiplicity of fiber products and idealization rings

### v2

@@ 3. Hilbert-Kunz Multiplicity of Fiber Product Rings
In this section, the rings $(R,\mathfrak{m}_{R},k)$, $(S,\mathfrak{m}_{S},k)$ and $(T,\mathfrak{m}_{T},k)$ will be considered Noetherian local rings as in the Definition 2.2. Also, assume that $R$ and $S$ have positive characteristic $p$ and $q$, respectively. In this case, the fiber product ring $R\times_{T}S$ is also a Noetherian local ring of positive characteristic $\max\{p,q\}$. Here, fiber product rings $R\times_{T}S$ will be considered under these conditions. To not be confused, note that in this section the rings $R$, $S$, and $T$ will also be considered as $R\times_{T}S$-modules.
3.1 Proposition.
@@
Next corollary provides a formula for the Hilbert-Kunz multiplicity of a fiber product $R\times_{T}S$ in terms of the Hilbert-Kunz multiplicities of $R$, $S$ and $T$, when $R$, $S$ and $T$ have the same characteristic.
3.2 Corollary.
@@
As an immediate consequence of Proposition 3.1, we obtain the formulas
@@
$e_{HK}(\mathfrak{m}_{A},B)=\lim_{e\to\infty}\frac{\ell_{A}\left(B/\mathfrak{m}^{[q]}_{A}B\right)}{q^{d}}=\lim_{e\to\infty}\frac{\ell_{B}\left(B/\mathfrak{m}_{B}^{[q]}\right)}{q^{d}}=e_{HK}(B).$
Therefore, by utilizing the formula for $\dim R\times_{T}S$ provided in Fact 2.5, the surjective ring homomorphisms of the diagram (2.5), and equality (3.4), we obtain the desired formulas. ∎
@@
Watanabe and Yoshida asked whether it is possible to determine the infimum of Hilbert-Kunz multiplicities over all unmixed Noetherian local rings of characteristic $p>0$ and dimension $d$ that are not regular. Due to this and the Watanabe-Yoshida conjecture (see Conjecture 5.3), some authors (see, e.g., [24, 6, 4, 1]) have determined lower and upper bounds for some unmixed local rings. However, it is also natural to ask whether it is possible to find lower and upper bounds for the Hilbert-Kunz multiplicity of Noetherian local rings that are not necessarily unmixed. In this sense, by using existing results, Corollary 3.2 helps us to find bounds for the fiber product ring (which is not always unmixed) by knowing the bounds of the components of the fiber product ring. Below, we provide one example among many to illustrate this fact:
@@
The inequality follows easily from Corollary 3.2 and [1, Theorem 4.12].
@@
We can use Corollary 3.2 to obtain a formula for the Hilbert-Kunz Multiplicity of the amalgamated duplication along an ideal.
@@
It follows directly by Corollary 3.2. ∎
@@
It is well known that regular local rings have Hilbert-Kunz multiplicity equal to one. In all the above-mentioned cases, using Corollary 3.2, we obtain $e_{HK}(R\times_{T}S)>1$ and then $R\times_{T}S$ can not be regular. ∎
@@
    By Corollary 3.2, we have $e_{HK}(R\times_{T}S)=e_{HK}(R)+e_{HK}(S)-e_{HK}(T)$. Also, we observe that as $R$, $S$ and $T$ are two-dimensional Cohen-Macaulay local rings, then the same is true for $R\times_{T}S$, by [2, Proposition 1.7]. For two-dimensional Cohen-Macaulay local rings, regularity is equivalent to unitary Hilbert-Kunz multiplicity, by [23, Theorem 2.5]). Then $R\times_{T}S$ is regular if and only if
@@
    The proof follows by Corollary 3.2 and [22, Theorem 1.5].
@@
To end this section, we will use the formulas given in Corollary 3.2 to calculate the Hilbert-Kunz multiplicity of some fiber product rings.
@@
Let $k$ be a field of characteristic $p>0$. If $R=k[[X_{1},\dots,X_{n}]]$ and $S=k[[Y_{1},\dots,Y_{m}]]$, then $R\times_{k}S\cong k[[\underline{X},\underline{Y}]]/(X_{i}Y_{j}:1\leq i\leq n,1\leq j\leq m)$, by Fact 2.6. Therefore, if $n>m$, then $\dim R>\dim S>\dim k$ and we have
@@
by Corollary 3.2. Note that (3.5) provides examples of non-regular local rings with Hilbert-Kunz multiplicity $1$. Also, if $n=m$, then $\dim R=\dim S>\dim k$ and we have
@@
for all positive integer $n\geq 4$, by Corollary 3.2.
@@
Let $k$ be a field of characteristic $p>0$. Let $R=k[[X]]$, $S=k[[Y]]$ and $T=K[[Z]]/(Z^{n})$, where $n\in\mathbb{N}$ and $n\geq 2$. Then $R\times_{T}S\cong k[[X,Y]](X^{n}Y-Y^{2})$, by [5, Corollary 4.2.7], taking $A=k[[X,Y]](X^{n}Y-Y^{2})$, $\mathfrak{p}=(X^{n}-Y)$ and $\mathfrak{q}=(Y)$. In this case, $\dim R=\dim S>\dim T$ and we have
@@
for all positive integer $n\geq 2$, by Corollary 3.2.
@@ 4. Hilbert-Kunz multiplicity of idealization rings
On other hand, the generators of the $R$-module $I^{[q]}(R\ltimes M)$ are of the form $r(s,m)$, where $r\in I^{[q]}$ and $(s,m)\in R\ltimes M$. Writing $r=\sum_{i=1}^{n}y_{i}x_{i}^{q}$, where $y_{i}\in R$ and $x_{i}\in I$, for $i=1,\dots,n$, we get
@@
4.5 Corollary.
  • (1)
    $e_{HK}(J,R\ltimes M)\leq(1+\mu(M))e_{HK}(I,R)$.
  • (2)
    If $\operatorname{pd}_{R}M=n<\infty$, then $e_{HK}(J,R\ltimes M)=\left(\sum_{i=0}^{n}(-1)^{i}\beta_{i}^{R}(M)+1\right)e_{HK}(I,R)$, where $\beta_{i}^{R}(M)$ denotes the i-th Betti number of $M$. In particular, if $M$ is a free $R$-module of rank $r$, then $e_{HK}(J,R\ltimes M)=(r+1)e_{HK}(I,R)$.
  • (3)
    If $R$ is a local domain, then $e_{HK}(J,R\ltimes M)=(\operatorname{rk}_{R}(M)+1)e_{HK}(I,R)$. Also, if we assume that $M$ is a torsion-free $R$-module of rank $r$, then $e_{HK}(J,R\ltimes M)=e_{HK}(I,M)+re_{HK}(R)$.
Proof.
  • (1)
    Considering a short exact sequence
    we get
  • (2)
    Taking a minimal free-resolution of $M$
    and using the additivity of Hilbert-Kunz multiplicity on short exact sequences we obtain
    $e_{HK}(I,M)=\sum_{i=0}^{n}(-1)^{i}\beta_{i}^{R}(M)e_{HK}(I,R).$
    Utilizing Theorem 2.9 we obtain the desired equality.
  • (3)
    The proof follows by Theorem 2.9 and itens (5),(6) of Proposition 2.1.
4.6 Example.
Indeed, if $M=R^{r}$, then $R\ltimes M\cong R[X_{1},\dots,X_{r}]/(X_{1},\dots,X_{r})^{2}$, by [3, Proposition 2.2], and the equality follows by the above corollary.
The next corollary provides some important structural information, in the sense of studying the regularity of local rings.
@@
    Using Corollary 4.3, we get $e_{HK}(R)=1$ and $e_{HK}(\mathfrak{m},M)=0$. As $R$ is unmixed it must be regular, by [22, Theorem 1.5]. Also, we have $\dim M<\dim R$, by item (3) of Proposition 2.1.
@@ 5. Some bounds for the Hilbert-Kunz multiplicity
Recently, Cox and Aberbach [8] proved that the Watanabe-Yoshida conjecture holds up to dimension $7$. Also, the conjecture is known in full generality for complete intersections by the work of Enescu and Shimomoto [11]
In our work, we calculated their Hilbert-Kunz multiplicity and gave conditions for fiber products and idealization rings not to be regular. The next results show new classes of $d$-dimensional Noetherian local rings of characteristic $p>0$ whose satisfy the part (1) of Conjecture of Watanabe and Yoshida by using the formulas maded in the sections 3 and 4.
@@
The proof follows by Corollary 3.2 and [11, Theorem 4.6]. ∎
@@
The proof follows by Corollary 4.3 and [11, Theorem 4.6]. ∎
@@
Let $(R,\mathfrak{m}_{R},k)$ and $(S,\mathfrak{m}_{S},k)$ be Noetherian local rings and suppose $\dim R,\dim S\geq 2$. Then $R\times_{k}S$ is not Cohen-Macaulay, by Fact 2.5 and [7, Remark 3.2]. It works event if $R$ and $S$ are regular local rings.

### v3

@@ 2. Setup and Basic facts
  • (10)
    If $(A,\mathfrak{m})\to(B,\mathfrak{n})$ is a flat local ring homomorphism such that $B/\mathfrak{m}B$ is a field, then $e_{HK}(I)=e_{HK}(IB)$.
@@ 3. Hilbert-Kunz Multiplicity of Fiber Product Rings
In this section, the rings $(R,\mathfrak{m}_{R},k)$, $(S,\mathfrak{m}_{S},k)$ and $(T,\mathfrak{m}_{T},k)$ will be considered Noetherian local rings as in the Definition 2.2. Also, assume that $R$ and $S$ have positive characteristic $p$ and $q$, respectively. In this case, the fiber product ring $R\times_{T}S$ is also a Noetherian local ring of positive characteristic $\max\{p,q\}$. During this section, fiber product rings $R\times_{T}S$ will be considered under these conditions. To not be confused, note that in this section the rings $R$, $S$, and $T$ will also be considered as $R\times_{T}S$-modules.
3.1 Lemma.
@@
Next proposition provides a formula for the Hilbert-Kunz multiplicity of a fiber product $R\times_{T}S$ in terms of the Hilbert-Kunz multiplicities of $R$, $S$ and $T$, when $R$, $S$ and $T$ have the same characteristic.
3.2 Proposition.
@@
As an immediate consequence of Lemma 3.1, we obtain the formulas
@@
$e_{HK}(\mathfrak{m}_{A},B)$ $=\lim_{e\to\infty}\frac{\ell_{A}\left(B/\mathfrak{m}^{[q]}_{A}B\right)}{q^{d}}$
$=\lim_{e\to\infty}\frac{\ell_{B}\left(B/\mathfrak{m}_{B}^{[q]}\right)}{q^{d}}$
$=e_{HK}(B).$
Therefore, by utilizing the formula for $\dim R\times_{T}S$ provided in Fact 2.5, the surjective ring homomorphisms of the diagram (2.5), and equality (3.4), we obtain the desired formulas, by observing that we need only to consider the components of the fiber product $R\times_{T}S$ whose have dimension equal to $\dim R\times_{T}S$. ∎
@@
Watanabe and Yoshida asked whether it is possible to determine the infimum of Hilbert-Kunz multiplicities over all unmixed Noetherian local rings of characteristic $p>0$ and dimension $d$ that are not regular. Due to this and the Watanabe-Yoshida conjecture (see Conjecture 5.3), some authors (see, e.g., [25, 6, 4, 1]) have determined lower and upper bounds for some unmixed local rings. However, it is also natural to ask whether it is possible to find lower and upper bounds for the Hilbert-Kunz multiplicity of Noetherian local rings that are not necessarily unmixed. In this sense, by using existing results, Proposition 3.2 helps us to find bounds for the fiber product ring (which is not always unmixed) by knowing the bounds of the components of the fiber product ring. Below, we provide one example among many to illustrate this fact:
@@
The inequality follows easily from Proposition 3.2 and [1, Theorem 4.12].
@@
We can use Proposition 3.2 to obtain a formula for the Hilbert-Kunz Multiplicity of the amalgamated duplication along an ideal.
@@
It follows directly by Proposition 3.2. ∎
@@
It is well known that regular local rings have Hilbert-Kunz multiplicity equal to one. In all the above-mentioned cases, using Proposition 3.2, we obtain $e_{HK}(R\times_{T}S)>1$ and then $R\times_{T}S$ can not be regular. ∎
@@
    By Proposition 3.2, we have $e_{HK}(R\times_{T}S)=e_{HK}(R)+e_{HK}(S)-e_{HK}(T)$. Also, we observe that as $R$, $S$ and $T$ are two-dimensional Cohen-Macaulay local rings, then the same is true for $R\times_{T}S$, by [2, Proposition 1.7]. For two-dimensional Cohen-Macaulay local rings, regularity is equivalent to unitary Hilbert-Kunz multiplicity, by [24, Theorem 2.5]). Then $R\times_{T}S$ is regular if and only if
@@
    The proof follows by Proposition 3.2 and [23, Theorem 1.5].
@@
To end this section, we will use the formulas given in Proposition 3.2 to calculate the Hilbert-Kunz multiplicity of some fiber product rings.
@@
Let $k$ be a field of characteristic $p>0$. If $R=k[[X_{1},\dots,X_{n}]]$ and $S=k[[Y_{1},\dots,Y_{m}]]$, then
$R\times_{k}S\cong\frac{k[[\underline{X},\underline{Y}]]}{(X_{i}Y_{j}:1\leq i\leq n,1\leq j\leq m)},$
by Fact 2.6. Therefore, if $n>m$, then $\dim R>\dim S>\dim k$ and we have
@@
by Proposition 3.2. Note that (3.5) provides examples of non-regular local rings with Hilbert-Kunz multiplicity $1$. Also, if $n=m$, then $\dim R=\dim S>\dim k$ and we have
@@
for all positive integer $n\geq 4$, by Proposition 3.2.
@@
Let $k$ be a field of characteristic $p>0$. Let $R=k[[X]]$, $S=k[[Y]]$ and $T=K[[Z]]/(Z^{n})$, where $n\in\mathbb{N}$ and $n\geq 2$. Then
$R\times_{T}S\cong\frac{k[[X,Y]]}{(X^{n}Y-Y^{2})},$
by [5, Corollary 4.2.7], taking $A=k[[X,Y]]/(X^{n}Y-Y^{2})$, $\mathfrak{p}=(X^{n}-Y)$ and $\mathfrak{q}=(Y)$. In this case, $\dim R=\dim S>\dim T$ and we have
@@
for all positive integer $n\geq 2$, by Proposition 3.2.
@@ 4. Hilbert-Kunz multiplicity of idealization rings
Then $J^{[q]}\subseteq I^{[q]}(R\ltimes M)$. On other hand, the generators of the $R$-module $I^{[q]}(R\ltimes M)$ are of the form $r(s,m)$, where $r\in I^{[q]}$ and $(s,m)\in R\ltimes M$. Writing $r=\sum_{i=1}^{n}y_{i}x_{i}^{q}$, where $y_{i}\in R$ and $x_{i}\in I$, for $i=1,\dots,n$, we get
@@
    Using Corollary 4.3, we get $e_{HK}(R)=1$ and $e_{HK}(\mathfrak{m},M)=0$. As $R$ is unmixed it must be regular, by [23, Theorem 1.5]. Also, we have $\dim M<\dim R$, by item (3) of Proposition 2.1.
4.6 Corollary.
  • (1)
    $e_{HK}(J,R\ltimes M)\leq(1+\mu(M))e_{HK}(I,R)$.
  • (2)
    If $\operatorname{pd}_{R}M=n<\infty$, then
    $e_{HK}(J,R\ltimes M)=\left(\sum_{i=0}^{n}(-1)^{i}\beta_{i}^{R}(M)+1\right)e_{HK}(I,R),$
    where $\beta_{i}^{R}(M)$ denotes the i-th Betti number of $M$. In particular, if $M$ is a free $R$-module of rank $r$, then $e_{HK}(J,R\ltimes M)=(r+1)e_{HK}(I,R)$.
  • (3)
    If $R$ is a local domain, then
    $e_{HK}(J,R\ltimes M)=(\operatorname{rk}_{R}(M)+1)e_{HK}(I,R),$
    Also, if in addition we assume that $M$ is a torsion-free $R$-module of rank $r$, then $e_{HK}(J,R\ltimes M)=e_{HK}(I,M)+re_{HK}(R)$.
Proof.
  • (1)
    Considering a short exact sequence
    we get
  • (2)
    Taking a minimal free-resolution of $M$
    and using the additivity of Hilbert-Kunz multiplicity on short exact sequences of finite $R$-modules, we obtain
    $e_{HK}(I,M)=\sum_{i=0}^{n}(-1)^{i}\beta_{i}^{R}(M)e_{HK}(I,R).$
    Utilizing Theorem 2.9 we obtain the desired equality.
  • (3)
    The proof follows by Theorem 2.9 and itens (5),(6) of Proposition 2.1.
4.7 Corollary.
Proof.
If $M=R^{r}$, then $R\ltimes M$ is naturally isomorphic to $R[X_{1},\dots,X_{r}]/(X_{1},\dots,X_{r})^{2}$, by [3, Proposition 2.2]. Therefore, the equality follows by the above corollary. ∎
@@ 5. Some bounds for the Hilbert-Kunz multiplicity
Recently, Cox and Aberbach [8] proved that the Watanabe-Yoshida conjecture holds up to dimension $7$. The first inequality of part (1) of the conjecture is known in full generality for complete intersections by the work of Enescu and Shimomoto [11]. Also, recently, Trivedi [22] proved the second inequality of part (1) of the conjecture, for $d\geq 4$ and $p>d-3$.
In our work, we calculated their Hilbert-Kunz multiplicity and gave conditions for fiber products and idealization rings not to be regular. The next results show new classes of $d$-dimensional Noetherian local rings of characteristic $p>0$ whose satisfy the part (1) of Conjecture of Watanabe and Yoshida by using the formulas presented in the sections 3 and 4.
@@
If in addition $d\geq 4$ and $p>d-3$, then $e_{HK}(R\times_{T}S)\geq e_{HK}(R_{d,p})\geq 1+m_{d}$ and item (1) of Conjecture 5.3 holds for the fiber product $R\times_{T}S$.
@@
The proof follows by Proposition 3.2, [11, Theorem 4.6] and [22, Theorem (A)]. ∎
@@
If in addition $d\geq 4$ and $p>d-3$, then $e_{HK}(R\bowtie I)\geq e_{HK}(R_{d,p})\geq 1+m_{d}$ and item (1) of Conjecture 5.3 holds for the amalgameted duplication $R\bowtie I$.
@@
If in addition $d\geq 4$ and $p>d-3$, then $e_{HK}(R\ltimes M)\geq e_{HK}(R_{d,p})\geq 1+m_{d}$ and item (1) of Conjecture 5.3 holds for the idealization ring $R\ltimes M$.
@@
The proof follows by item (1) of Corollary 4.3, [11, Theorem 4.6] and [22, Theorem (A)]
@@
Let $(R,\mathfrak{m}_{R},k)$ and $(S,\mathfrak{m}_{S},k)$ be Noetherian local rings and suppose $\dim R,\dim S\geq 2$. Then $R\times_{k}S$ is not Cohen-Macaulay, by Fact 2.5 and [7, Remark 3.2].
Competing interests: The authors declare none.